OpenStudy (anonymous):

(2n^4h^3)^4 I don't even....

hero (hero):

How'd you get from adding fractions to this?

OpenStudy (anonymous):

Because he is posting problems from a Connections Academy Test.

hero (hero):

Yes, that pretty much explains it

OpenStudy (anonymous):

(2 n^4 h^3)^4 = 2^4 n^(4 * 4) h^(3 * 4) = 16 n^16 h^12

OpenStudy (anonymous):

I hope that helps :)